OpenStudy (anonymous):

Please help! I don't understand how to do this: Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = (x-7)/(x+3) and g(x) = (-3x-7)/x-1

OpenStudy (phi):

f(g(x)) means start with f(x)= (x-7)/(x+3) everywhere you see an x, replace it with g(x). It gets messy: \[ f(g(x))= \frac{ \frac{-3x-7}{(x-1)} - 7}{ \frac{-3x-7}{(x-1)}+3} \] now simplify. One good idea is to start by multiplying the top and bottom by (x-1)

OpenStudy (anonymous):

so if i multiply the top and bottom by x-1 would i multiply the 7 and 3 by it as well?

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

the advantage to it, is you get rid of the denominators atop and below

OpenStudy (anonymous):

i got -4x/10 i dont see how that proves it though

OpenStudy (phi):

you goofed somewheres.... after multiplying the top by x-1 what did you get for the new top?

OpenStudy (phi):

you are working with this, right ? \[ f(g(x))= \frac{ \frac{-3x-7}{(x-1)} - 7}{ \frac{-3x-7}{(x-1)}+3} \]

OpenStudy (jdoe0001):

one sec

OpenStudy (anonymous):

yes i got 3x-7-7x+7 and then i simplified

OpenStudy (jdoe0001):

did you this after you multiplied as phi suggested $$ \cfrac{-3x-7-7(x-1)}{-3x-7+3(x-1)} $$ ?

OpenStudy (jdoe0001):

sorry I'm a bit lagged :/

OpenStudy (jdoe0001):

did you get that I meant

OpenStudy (phi):

it's -3x -7 -7x +7

OpenStudy (anonymous):

ok now i got -x, is that right?

OpenStudy (phi):

closer, what did you get for the bottom?

OpenStudy (jdoe0001):

ahemm, (g(x)) = x and g(f(x)) = x.

OpenStudy (anonymous):

just kidding i got x. -3x-7+3x-3

OpenStudy (phi):

yes -10x/(-10) = x now do the same thing for g(f(x))

OpenStudy (anonymous):

would that start out like this?: ((7x-1/x+3)-3x-7)/((7x-1/x+3) + x-1 ?

OpenStudy (phi):

no, you replace all x's in g(x) with f(x) g(x) = (-3x-7)/(x-1) wherever you see an x, erase it, and squeeze in f(x)

OpenStudy (anonymous):

got it thank you so much!